Adams Camps, Inc. leases the land on which it builds camp sites. Adams is considering opening a new site on land that requires $2,200 of rental payment per month. The variable cost of providing service is expected to be $5 per camper. The following chart shows the number of campers Adams expects for the first year of operation of the new site:

Jan. Feb. Mar. Apr. May June July Aug. Sept. Oct. Nov. Dec. Total
170 270 230 220 360 520 670 670 370 400 200 320 4,400


Required

Assuming that Adams wants to earn $8 per camper, determine the price it should charge for a camp site in February and August. (Do not round intermediate calculations.)
Answer #1

Calculate price should charge :

February August
Fixed cost 2200 2200
Variable cost 270*5 = 1350 670*5 = 3350
Desired profit 270*8 = 2160 670*8 = 5360
Total 5710 10910
No of camper 270 670
Price charge 21.15 16.28